OpenStudy (anonymous):

An apple orchard is selling apple cider is spherical sipping containers. If each container is 7cm in radius, what volume of liquid is in each container ?

HanAkoSolo (jamierox4ev3r):

the formula for the volume of a cylinder is: \(V=bh\)

HanAkoSolo (jamierox4ev3r):

wait are you asking is for a sphere! oh im sorry :P

HanAkoSolo (jamierox4ev3r):

the formula for the volume of a sphere is \[V=\frac{ 4 }{ 3 }\pi*R^3\]

HanAkoSolo (jamierox4ev3r):

\(7^3\)=343

HanAkoSolo (jamierox4ev3r):

343*3.14=1077.02

HanAkoSolo (jamierox4ev3r):

ans *4/3=1436.026667

HanAkoSolo (jamierox4ev3r):

the volume, rounded to the nearest cm, is 1,436cm^3

HanAkoSolo (jamierox4ev3r):

@xoxox123 does this make sense?

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

Is the answer\[\frac{ 1372 \pi }{ 3 }\]?

OpenStudy (anonymous):

its supposed to be a number .

OpenStudy (anonymous):

To what decimal place is It rounded too?

OpenStudy (anonymous):

You don't round unless the problem says too. In this case it doesn't

OpenStudy (anonymous):

Than try 1437 cm ^3.

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

did you put the units cm^3 for both trys?

OpenStudy (anonymous):

are you supposed to write it like this 1437 cm^3